OpenStudy (anonymous):

at 0.500 atm and 15C a sample of gas occupies 120 L what volume does it occupy at 0.250 atm and 10C

OpenStudy (anonymous):

Convert your temperature to unit Kelvin, then use the combined gas law \[\frac{ P_1V_1 }{ T_1 } = \frac{ P_2V_2 }{ T_2 }\]

OpenStudy (anonymous):

0.500(120)/288= 0.250(?)/283

OpenStudy (anonymous):

yup